Subtract 1 quarter from 5
Answer by oberobic(2304)

Convert 5 to quarters.
x/4 = 5
x = 20
.
20/4 - 1/4 = 19/4
.
19/4 = 16/4 + 3/4 = 4 + 3/4, which is usually written as 4 3/4.
.
Of course, this work could have been done using decimals.
1/4 = 0.25
5 - 0.25 = 4.75
4.75 = 4 3/4
.
Done.